Use the function f(x)=2^x/x to answer the question.

Which combination could result in the function f(x)?


  • adding a quadratic function and a rational function
  • dividing an exponential function by a linear function
  • adding an exponential function and a linear function


(select the correct option)



1
Expert's answer
2021-06-22T07:12:23-0400

Let us use the function "f(x)=\frac{2^x}{x}" to answer the question: which combination could result in the function "f(x)"? Taking into account that the function "g(x)=2^x" is an exponential function and the function "h(x)=x" is a linear function, we conclude that the function "f(x)=\frac{g(x)}{h(x)}=\frac{2^x}{x}"is a result of dividing an exponential function by a linear function.


Answer: dividing an exponential function by a linear function